[The following information applies to the questions displayed below.]

The equity sections for Atticus Group at the beginning of the year (January 1) and end of the year (December 31) follow.

Stockholders’ Equity (January 1)
Common stock—$4 par value, 100,000 shares
authorized, 30,000 shares issued and outstanding
$ 120,000
Paid-in capital in excess of par value, common stock 80,000
Retained earnings 320,000
Total stockholders’ equity $ 520,000
Stockholders’ Equity (December 31)
Common stock—$4 par value, 100,000 shares
authorized, 35,400 shares issued, 3,000 shares in treasury
$ 141,600
Paid-in capital in excess of par value, common stock 112,400
Retained earnings ($30,000 restricted by treasury stock) 420,000
674,000
Less cost of treasury stock (30,000 )
Total stockholders’ equity $ 644,000


The following transactions and events affected its equity during the year.

Jan. 5 Declared a $0.40 per share cash dividend, date of record January 10.
Mar. 20 Purchased treasury stock for cash.
Apr. 5 Declared a $0.40 per share cash dividend, date of record April 10.
July 5 Declared a $0.40 per share cash dividend, date of record July 10.
July 31 Declared a 20% stock dividend when the stock’s market value was $10 per share.
Aug. 14 Issued the stock dividend that was declared on July 31.
Oct. 5 Declared a $0.40 per share cash dividend, date of record October 10.

5. How much net income did the company earn this year

The answer is not 156880
Answer #1

Cash dividend = (30000*.40+27000*.40+27000*.40+32400*.40) = 46560

Stock dividend = 27000*20%*10 = 54000

Net income = Beginning equity-Cash dividend-stock dividend-Ending equity

= 320000-46560-54000-420000

Net income = 200560